OpenStudy (anonymous):

9e^5x/5x-3 find derivative

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=9e^{5x}, f'(x)=45e^{5x}, g(x)=5x-3, f'(x)=5\]

OpenStudy (anonymous):

i don't know, i didn't compute it, but that is what you need above if you only want the answer, try this http://www.wolframalpha.com/input/?i=%289e^5x%29%2F%285x-3+%29

OpenStudy (agent0smith):

@lexi69 the point of this site is not just to give you the answer...

OpenStudy (goformit100):

compute it @lexi69

OpenStudy (anonymous):

agent0smith why are you on my question if you not going to respond to it get lost who said i was on here just to get all the answers im trying learn it too

OpenStudy (anonymous):

@lexi69 I heard someone asked "whats the answer", aren't you the one???

OpenStudy (agent0smith):

@lexi69 I'm on your question to tell you that the point of the site is to help you understand how to find the answer on your own.